### v1

# Understanding the Skill Gap in Recurrent Language Models: The Role of the Gather-and-Aggregate Mechanism

Abstract
State-space models (SSMs) have recently emerged as efficient recurrent alternatives to Transformers for processing long sequences with fixed state sizes. However, this efficiency comes at the cost of reduced capabilities in certain algorithmic tasks, such as retrieval of past context. In this work, we examine how such in-context retrieval operates within Transformer- and SSM-based language models. We find that both architectures develop the same fundamental Gather-and-Aggregate (G&A) mechanism. A Gather Head first identifies and extracts relevant information from the context, which an Aggregate Head then integrates into a final representation. Across both model types, G&A concentrates in just a few heads, making them critical bottlenecks even for benchmarks that require a basic form of retrieval. For example, we show that disabling a single Gather or Aggregate Head of a pruned Llama-3.1-8B severely degrades its ability to retrieve the correct answer letter in MMLU, reducing its accuracy from 66% to 25% (random guessing). Moreover, this finding suggests that in-context retrieval can obscure the limited knowledge demands of certain tasks. Although the pruned model performs well on MMLU with functional retrieval, it struggles on other knowledge benchmarks. This G&A bottleneck extends to GSM8K, BBH, and dialogue tasks—all areas where SSMs typically underperform. Given the significance of G&A capability in performance, we show that the known retrieval challenges in SSMs manifest in how they implement G&A within these heads, leading to smoother attention patterns rather than the sharp token transitions that effective G&A relies on. Thus, while a gap exists between Transformers and SSMs in implementing in-context retrieval, it is limited to just a few heads, rather than a property of the entire language model. This insight points to a possible unified explanation for empirical performance differences between Transformers and SSMs while also highlighting ways to effectively combine their strengths. For example, we show that in pretrained hybrid models, where SSMs are interspersed with a few attention layers, the attention components naturally take on the role of Aggregate Heads. Similarly, in a pretrained pure SSM, replacing a single G&A head with an attention-based variant significantly improves retrieval capabilities, leading to higher benchmark scores.
1 Introduction
The rise of Transformers has reshaped modern language modeling, driving breakthroughs in natural language processing. Yet, Transformers’ quadratic scaling with sequence length has renewed interest in recurrent alternatives, particularly state-space models (SSMs). SSMs offer linear scaling with a fixed-size memory while remaining competitive across many tasks. However, replacing a large, explicit history cache with a compact state representation poses a fundamental tradeoff: SSM-based language models often struggle with the skill of “in-context retrieval”—the ability to precisely locate and extract tokens presented earlier in the sequence, such as recalling a phone number mentioned several paragraphs ago (Jelassi et al., 2024). This “algorithmic” challenge, which Transformers handle more effectively through direct attention to past tokens, has been identified as a key factor in the performance gap between Transformer models and SSM variants (Arora et al., 2023).
While previous work has explored retrieval limitations in SSMs (Wen et al., 2024; Jelassi et al., 2024; Waleffe et al., 2024), these studies have either treated language models as black boxes or focused on simplified theoretical setups. Meanwhile, research on Transformers has shown that algorithmic capabilities, including retrieval, are often concentrated in specific attention heads (Lieberum et al., 2023; Wu et al., 2024). This raises a key question: Can the performance gap between Transformer- and SSM-based language models be traced to a few algorithmic capabilities within a small subset of heads? Clarifying this will enable a more targeted comparison of the two architectures, helping determine whether SSMs can serve as viable building blocks for language modeling or require fundamental modifications.
The MMLU benchmark (Hendrycks et al., 2021) provides a particularly sharp lens into this performance gap. While SSMs can match Transformer results on many knowledge benchmarks, they tend to perform worse on MMLU unless trained significantly longer (Waleffe et al., 2024). This gap may stem from MMLU’s evaluation format: rather than scoring the likelihood of full-text answer candidates, models are required to select a letter label (A, B, C, or D) corresponding to the correct option. This format introduces an additional layer of complexity—models must not only possess the knowledge but also retrieve the corresponding letter from the context, creating a dual challenge of both knowledge and retrieval (Lieberum et al., 2023). To separate these distinct challenges, we analyze how each architecture handles the task. We find a clear division of roles: while factual knowledge is extracted from the model’s weights in a distributed manner across layers and heads, both Transformer- and SSM-based models rely on a small number of heads to retrieve the correct label from the input context.
(a) Gather Head
(b) Aggregate Head
[…]
This sharp separation between knowledge and in-context retrieval underscores that it is not knowledge alone, but the skill of retrieving it from context, that drives success on benchmarks where SSMs typically fall short. Disabling retrieval heads leads to steep performance drops across architectures on tasks that require accessing and reusing prior information—such as mathematical reasoning (GSM8K), multi-step logical reasoning (BBH), and dialogue formats (ARC-C Chat)—while other benchmarks remain largely unaffected. This effect is particularly evident in MMLU, where in-context retrieval capabilities appear to be more important than stored knowledge. For example, our experiments with a pruned 4B-parameter Llama-3.1-8B submodel show that despite performing poorly on most benchmarks, it still achieves a strong MMLU score of 66%. However, when we disable just a single retrieval head, MMLU accuracy plummets to 25%, approximately the level of random guessing. This dramatic decline reveals that MMLU success depends more on retrieval capabilities—where Transformers maintain a clear advantage over SSMs—than on stored knowledge.
The critical impact of in-context retrieval motivates our investigation of the underlying mechanism of these heads. We find that both Transformer- and SSM-based language models converge on a similar retrieval framework: a Gather-and-Aggregate (G&A) mechanism. Specifically, a Gather Head identifies and condenses segments of input tokens into a single vector representation (Figure 1(a)), which an Aggregate Head subsequently processes to extract the necessary information (Figure 1(b)). This finding indicates that, despite architectural differences, the two language models rely on similar retrieval strategies rather than fundamentally different approaches. Consequently, their retrieval gap reduces to the effectiveness with which a few specific heads implement the G&A mechanism.
Building on this observation, we compare the G&A implementation of the two language models. As expected, we find that the known retrieval challenges of SSMs manifest in language models through the way a few key heads implement G&A. SSMs exhibit smoother attention patterns, making it difficult to execute the sharp token shifts needed for G&A, requiring more of their instances to achieve comparable performance. Moreover, we show how hybrid models (Glorioso et al., 2024; Lieber et al., 2024), which integrate a few attention layers into predominantly SSM-based architectures, overcome SSM’s limitations. When trained from scratch, hybrid models naturally delegate Aggregate Head functionality to attention layers. Conversely, in pretrained SSM models, replacing a single G&A instance with an attention-based one yields a significantly greater MMLU improvement than any other attention placement. While these results suggest that attention has a higher capacity for implementing G&A, they simultaneously highlight the mechanism’s concentrated nature. This indicates that not every layer needs to cache the full history—paving the way for more efficient architectures that store past tokens in fewer components.
Overall, our contributions are as follows.
  1. 1.
    Gather-and-Aggregate (G&A) Retrieval Framework. We formalize a unified mechanism—comprising Gather Heads and Aggregate Heads—that enables in-context retrieval, extending previous work by Lieberum et al. (2023); Wu et al. (2024) from Transformers to SSM-based language models and advancing mechanistic interpretability of SSMs. Our analysis reveals that G&A functionality is highly concentrated in SSMs as well, with just a few heads among thousands performing this function. Furthermore, we demonstrate that SSMs’ inherent retrieval limitations significantly impair their ability to implement G&A as effectively as Transformers.
  2. 2.
    Empirical Analysis of G&A across Benchmarks. We conduct extensive empirical analysis across Transformers, SSMs, and hybrid architectures, revealing how G&A heads critically determine performance on retrieval-intensive benchmarks such as MMLU, GSM8K, and BBH. Our findings suggest that the broader performance gap between Transformers and SSMs may stem from specific algorithmic limitations in a small subset of heads rather than general language modeling deficiencies. Most notably, we show that the difficulty of MMLU arises more from in-context retrieval demands than from knowledge representation, as models can perform well on MMLU despite struggling with other tasks, but fail catastrophically without functional G&A mechanism.
  3. 3.
    Understanding the Effectiveness of Hybrid Models: We demonstrate that hybrid models address the retrieval limitations of SSMs in implementing G&A. When trained from scratch, these models naturally assign Aggregate Head functionality to attention layers. Alternatively, “hybridizing” a pretrained SSM by replacing a single G&A component with an attention-based G&A significantly improves performance. Furthermore, we provide evidence that both architectures tend to develop these heads in the middle of the network, suggesting an effective placement for such hybrids.
2 Related Work
Sequence Models.
Sequence models, particularly Transformer-based language models with self-attention (Vaswani et al., 2023; Brown et al., 2020; Touvron et al., 2023), dominate modern language modeling due to their scalability and strong performance. Recently, recurrent models, like structured state-space models (SSMs) (Gu et al., 2022; Gu and Dao, 2023; Dao and Gu, 2024), have gained focus for efficiently processing long sequences. These models replace quadratic attention with sub-quadratic methods, offering competitive performance while maintaining efficiency. Other recurrent models with similar characteristics include (Dao et al., 2022; Sun et al., 2023; Katharopoulos et al., 2020; Yang et al., 2024a; Zhang et al., 2024b; Beck et al., 2024; Qin et al., 2024; Peng et al., 2024; Yang et al., 2024b). In this work, we use “SSM” to refer to these models as well, as our findings are broadly applicable.
Despite SSMs’ efficiency, they struggle with tasks requiring precise token interactions, such as in-context retrieval. Their fixed-size hidden state limits memory capacity, making it difficult to retain detailed past information. To address these limitations, recent work (Lieber et al., 2024; Glorioso et al., 2024; Waleffe et al., 2024; Ren et al., 2024; Wang et al., 2024a; Paliotta et al., 2025) has adopted a hybrid approach, integrating Attention layers with a majority of SSM layers, trading efficiency for improved capabilities such as retrieval.
Throughout this study, we examine three different architectures: Llama-3.1-8B-Instruct (Touvron et al., 2023), a standard Transformer that achieves high accuracy on various NLP benchmarks; Falcon-Mamba-7B-Instruct (Zuo et al., 2024), a Mamba-1 SSM model without MLP layers; and Llamba-8B-Untied (Bick et al., 2025), a Mamba-2 SSM model distilled from Llama-3.1-8B using MOHAWK (Bick et al., 2024), which alternates MLP layers with SSMs. We refer to these as Llama-3.1-8B, Llamba-8B, and Falcon-Mamba-7B, respectively. Furthermore, Section 5 examines Zamba-2-7B (Glorioso et al., 2024), a hybrid model where Mamba2 layers outnumber attention layers by a 6:1 ratio, with only two attention layers shared across the entire model.
Central to understanding the performance differences between these architectures is the concept of token mixing. Token mixing refers to how sequence models propagate and combine information across different token positions. In self-attention, each token selectively incorporates information from other tokens based on learned weights. SSMs, in contrast, mix tokens using structured recurrences that enforce a fixed computation pattern. Ali et al. (2024); Dao and Gu (2024) link SSM’s token-mixing to the matrix-multiplication form of self-attention. They showed that Mamba-1 mimics thousands of attention heads with a head dimension of 1, interacting with the input in a multi-value manner. Dao and Gu (2024) introduced Mamba-2, which uses fewer heads with larger dimensions (Dao and Gu, 2024; Bick et al., 2025; Waleffe et al., 2024), making each head more expressive than in Mamba-1. These insights enable us to view all mentioned sequence models as structured matrices operating along the time dimension. For that reason, we generalize token-mixing mechanisms beyond traditional self-attention, and use throughout this work the term Temporal Mixing Heads (or simply Mixing Heads).
Retrieval Tradeoffs Between Transformers and SSMs.
SSM-based language models rely on compressed state representations, where information is continually updated in a low-dimensional hidden state—unlike Transformers, which cache their full context. These inherent differences have been observed across multiple benchmarks. Waleffe et al. (2024) found that Mamba-based models require extensive training to perform well on MMLU, while Wen et al. (2024) identified weaknesses in in-context retrieval tasks, such as associative recall. Jelassi et al. (2024) theoretically and empirically confirmed that SSMs have difficulty with precise copying. Arora et al. (2023) further demonstrated this gap, finding that associative recall capabilities account for majority of the performance difference between attention-based and gated-convolution models. Park et al. (2024) benchmarked Mamba and Transformer models across a range of in-context learning tasks, showing that SSMs underperform on tasks requiring non-standard retrieval, and proposed a hybrid architecture to address this limitation. Arora et al. (2025) analyzed the recall-throughput tradeoff across architectures, showing that attention-based models occupy a different region of the tradeoff frontier than SSMs, with stronger recall but higher computational cost. Blouir et al. (2025) complemented this view by enhancing retrieval in SSMs through training improvements without architectural changes. Unlike previous work, our study reveals that while these limitations exist, they originate from a small subset of heads responsible for retrieval rather than the entire language model.
Mechanistic Interpretability of Mixing Heads.
Most interpretability studies focus on attention heads in LLMs (Elhage et al., 2021; Todd et al., 2024; Lieberum et al., 2023; Zhang et al., 2024a; Olsson et al., 2022; Yu et al., 2024; Merullo et al., 2024; Rai et al., 2024; Zheng et al., 2024; Wu et al., 2024; Tulchinskii et al., 2024; Sanford et al., 2024), while relatively little attention has been given to recurrent mechanisms (Ali et al., 2024). Olsson et al. (2022) and Elhage et al. (2021) showed that induction heads emerge to support in-context learning. Subsequent work traced their training dynamics (Bietti et al., 2023), and mapped factual recall circuits (Meng et al., 2023).
In a more focused study, Lieberum et al. (2023) analyzed Chinchilla 70B on multiple-choice tasks and identified two key attention heads: the Content Gatherer Head and the Correct Letter Head, which work in conjunction with MLPs to select correct labels. In addition, Wu et al. (2024) identified Retrieval Heads—a small subset of Transformer heads essential for retrieval that comprise less than 5% of all heads but whose removal significantly disrupts retrieval, chain-of-thought reasoning, and increases hallucinations.
Our contribution refines the understanding of their roles in supporting retrieval. We show that in both Transformers and SSMs, retrieval emerges not from a single head but through a coordinated mechanism: one head gathers the location of the target, and another aggregates the content. This framing extends prior work in three key ways. First, the “Correct Letter” heads identified by Lieberum et al. (2023) exhibit broader retrieval behavior than previously reported, going beyond multiple-choice selection. Second, retrieval depends on coordinated interactions between heads, challenging the interpretation in Wu et al. (2024) that treated individual heads as independently responsible. Third, we demonstrate that this mechanism also arises in SSMs—despite their lack of attention—highlighting its architectural generality. We introduce the term Aggregate Head to better reflect these empirical findings, while retaining “Gather” as shorthand for the “Content-Gatherer” heads described Lieberum et al. (2023).
Concurrent with our work, Wiegreffe et al. (2025) expanded on Lieberum et al. (2023)’s findings beyond MMLU and Chinchilla-70B, investigating model behavior throughout training and analyzing cases of poor performance. They created synthetic tasks to separate knowledge from multiple-choice formatting abilities and uncovered mechanisms for how models adapt to different answer choice symbols. While their objectives align with ours, our research specifically targets the “algorithmic” in-context retrieval aspect of multiple-choice formatting across various architectures, contrasting this with models’ underlying knowledge capabilities.
3 A Case Study: Retrieval in MMLU
Figure 2: Performance of Llama-3.1-8B on knowledge tasks and MMLU as layers are gradually pruned, starting from the last. The leftmost bar shows the full 32-layer model, while subsequent bars represent pruned versions. Key observations: (1) MMLU scores stay stable despite declining knowledge tasks performance; (2) a sharp MMLU drop identifies a critical skill-encoding layer. (3) 30% of the model’s original score on knowledge tasks is sufficient to reach a state-of-the-art MMLU score.
MMLU covers 57 tasks that include elementary mathematics, US history, computer science, law, and more. To achieve high accuracy in this test, models must have extensive world knowledge and problem solving ability (Hendrycks et al., 2021). MMLU covers 57 diverse tasks, including elementary mathematics, U.S. history, computer science, and law. Thus, models must possess extensive world knowledge to perform well (Hendrycks et al., 2021). In addition to the knowledge requirement, the MMLU evaluation format requires selecting a letter label (A-D) and retrieving it rather than scoring complete answers, demanding a basic form of retrieval ability.
___ is the central node of 802.11 wireless operations.
A. WPA
B. Access Point
C. WAP
D. Access Port
Answer:
SSMs tend to perform worse on MMLU unless trained significantly longer (Waleffe et al., 2024), despite matching Transformer performance on many knowledge benchmarks. Given known retrieval limitations of SSM architectures (Jelassi et al., 2024), we ask whether their performance gap with Transformers on MMLU stems from difficulties in retrieving rather than acquisition of knowledge?
This section presents two key findings:
  1. 1.
    Models can perform well on MMLU with limited knowledge, suggesting that, for the evaluated models, the task primarily reflects in-context retrieval. This adds to concerns raised by prior work, including Wang et al. (2024b); Zhao et al. (2024).
  2. 2.
    Both Transformer- and SSM-based models encode retrieval abilities in a small set of specialized heads, extending prior work that focused solely on Transformers (Lieberum et al., 2023; Wu et al., 2024).
These insights, combined with earlier findings about perplexity gaps between attention-based models and SSM variants on the Pile dataset (Arora et al., 2023), support our main thesis: the general performance gap between Transformer- and SSM-based language models can be traced to specific algorithmic capabilities within a small subset of heads.
Our investigation compares the attention-based Transformer, Llama-3.1-8B-Instruct, with two SSM-based architectures: Falcon-Mamba-7B-Instruct (based on Mamba-1) and Llamba-8B-untied (based on Mamba-2). For brevity, we omit instructional and embedding-tying descriptors throughout our analysis. We evaluate performance on MMLU alongside a suite we term Knowledge-Focused Tasks. These include ARC-Challenge, ARC-Easy, PIQA, Winogrande, OpenBookQA, and HellaSwag—tasks selected because they primarily assess factual knowledge and commonsense reasoning with minimal dependence on retrieval from recent context. Our analysis follows a bottom-up approach, identifying discrepancies between MMLU and other knowledge-focused tasks, then tracing them back to retrieval. Specifically, we demonstrate:
  • Section 3.1 outlines the benchmarks used in our analysis, distinguishing MMLU from knowledge-focused tasks based on their reliance on retrieval.
  • Section 3.2 shows that MMLU performance initially appears to depend on a specific critical layer rather than the network’s overall stored knowledge, based on cumulative layer pruning.
  • Section 3.3 builds on this finding through more targeted experiments, revealing that MMLU actually relies on the coordination between two essential layers, both of which must be present for successful performance.
  • Section 3.4 refines this perspective by shifting from layers to individual heads, revealing that MMLU disproportionately relies on a few specific heads rather than distributed knowledge.
  • Section 3.5 demonstrates these heads are part of a retrieval mechanism consistently present across all architectures.
Figure 3: Performance of Llama-3.1-8B-Minimal with layers removed and restored. We remove one layer at a time, evaluate performance on MMLU and knowledge tasks’ score, and then restore the layer before proceeding to the next. The results reveal that both layers 16 and 17 are critical for MMLU performance, as pruning either of the last two layers of the minimal model causes a significant drop on MMLU scores while causing only a minor drop in scores on knowledge-focused tasks.
3.1 Evaluation Benchmarks.
Throughout the paper, we evaluate performance on two categories of benchmarks: MMLU and what we term Knowledge-Focused Tasks. MMLU’s format distinctively presents multiple-choice questions across 57 diverse subjects, requiring models to process lengthy contexts containing both questions and answer choices, then identify the correct option by outputting the corresponding letter label (typically A, B, C, D). While MMLU is widely used as a benchmark, some researchers have raised questions about its evaluation approach (Alzahrani et al., 2024; Wang et al., 2024b; Zhao et al., 2024) For instance, Wang et al. (2024b) argue that MMLU performance may not necessarily indicate true comprehension by large language models, and studies by Zhao et al. (2024) have highlighted issues of benchmark contamination that inevitably affect evaluation integrity.
In contrast, our Knowledge-Focused Tasks—including ARC-Challenge and ARC-Easy (Clark et al., 2018), PIQA (Bisk et al., 2019), Winogrande (Sakaguchi et al., 2019), OpenBookQA (Mihaylov et al., 2018), and HellaSwag (Zellers et al., 2019)—were selected because they primarily assess factual knowledge and commonsense reasoning with minimal dependence on retrieval from recent context. The fundamental difference lies in what the models are asked to predict: while MMLU requires models to retrieve a letter label (A, B, C, or D) after processing question-answer pairs, these other benchmarks are evaluated by scoring the likelihood each model assigns to full candidate answers. In this setup, we retrieve the complete text of each answer option in advance and present it directly to the model. The model’s task is then to assign a likelihood score to each one, rather than selecting a label. This evaluation approach removes the burden of interpreting or extracting answer choices from the context and instead emphasizes how well the model’s internal knowledge supports each complete answer.
3.2 A Crucial Layer for MMLU
Motivated by the struggle of SSMs with MMLU (Waleffe et al., 2024), and by the fact that Transformers use specialized components to solve multiple-choice questions (Lieberum et al., 2023), we examine whether solving this task utilizes different components in SSM-based language models as well. We start our investigation by isolating specific layers’ contributions. Since all language models utilize residual skip connections, eliminating a layer could reveal its specific contribution, while keeping the rest of the information flowing.
We iteratively remove one layer at a time from the model’s end and measure performance on MMLU against knowledge-focused tasks. For Llama-3.1-8B and Llamba-8B, we remove entire layers, including both mixer and MLP components, since MLPs likely depend on mixer outputs. The experiments provided in Figure 9 show similar results when the MLPs are kept intact.
Figure 2 presents the performance trends for Llama-3.1-8B on both knowledge tasks and MMLU as layers are progressively removed, and Figure 8 extends this analysis to Falcon-Mamba-7B and Llamba-8B models. All architectures revealed a striking pattern: while performance on knowledge-focused benchmarks decreased uniformly as layers were pruned, MMLU performance remained remarkably stable until a specific layer was removed. At this point, MMLU performance dropped dramatically, suggesting this particular layer encodes crucial skills required for MMLU. This observation also raises an intriguing question about how these models can maintain high MMLU scores while apparently retaining only a fraction of their knowledge capacity.
3.3 MMLU Relies on Two Crucial Layers
Based on insights from Figure 2, we refer to the minimal model as the smallest subnetwork that retains all layers up to (and including) the layer before the sharp drop in MMLU performance. We now repeat the previous experiment, this time on the minimal model, with a slight modification. Instead of cumulatively removing one layer at a time from the end of the network, we remove a single layer, evaluate the model performance without it, and then restore the layer before proceeding to the next.
Figure 3 illustrates the results of this experiment in Llama-3.1-8B, and Figure 10 shows the results for Falcon-Mamba-7B and Llamba-8B as well. The findings reveal a surprising insight on the minimal model: not one, but two layers are critical for the performance of MMLU. Unlike earlier layers, the last two layers of the minimal model show no correlation between MMLU score and knowledge task performance. Pruning either layer causes a significant drop in MMLU accuracy while causing only a minor drop in scores on knowledge-focused tasks, underscoring their essential role in the task.
Heads Kept in a Layer Metrics (%)
0-15 16 17 MMLU Knowledge Tasks
0,1,…,31 22 24 66.32 39.09
0,1,…,31 $\emptyset$ 24 24.36 39.18
0,1,…,31 22 $\emptyset$ 25.59 39.21
0,1,…,31 $\emptyset$ $\emptyset$ 25.56 39.21
Table 1: Evaluating the Role of Critical Heads in Llama-3.1-8B-Minimal’s Performance. This table presents the impact of two critical attention heads (L16H22 and L17H25) on MMLU performance while keeping all preceding layers intact. Four configurations were tested: (1) both heads retained, (2) only the last-layer head retained, (3) only the second-to-last-layer head retained, and (4) neither head retained. Results show that these heads must work together to sustain high MMLU performance ($>$66%), while removing either or both drops accuracy to random guessing (25%).
3.4 Zooming In: MMLU Relies on Two Crucial Heads
Having established the importance of two specific layers in Section 3.3, we now refine our analysis from the layer level to the head level, investigating how individual attention heads within these critical layers contribute to each language model’s performance.
To precisely identify the crucial heads, we conducted a systematic ablation study on the minimal model defined in Section 3.3 (where only the last two layers are critical). We systematically zeroed out the outputs of individual heads in these two critical layers and measured the impact on both MMLU and knowledge-focused tasks. This analysis revealed a remarkably small subset of heads whose removal drastically reduced MMLU accuracy while leaving knowledge benchmark performance intact. We designated these as ”critical heads” - specifically, the heads whose removal significantly impacted MMLU scores.
To further validate the importance of these identified critical heads, we removed all noncritical heads from the last two layers and evaluated performance across four configurations: (1) no critical heads retained, (2) only critical heads from the last layer retained, (3) only critical heads from the second last layer retained, and (4) critical heads in both layers retained. The results, presented in Table 1 for Llama-3.1-8B and in Tables 6 and 5 for the other models, demonstrate that MMLU performance depends on just a few mixing heads rather than the model’s broader knowledge representation. Disabling these specific heads causes accuracy to plummet to random guessing levels. Remarkably, in Llama-3.1-8B, a single head is sufficient to achieve over 66% MMLU accuracy, while in Falcon-Mamba-8B, just four Mamba-1 channels reach 52.26% accuracy.
3.5 Retrieval Emerges as a Key Factor in MMLU
Our earlier findings demonstrated that two attention heads from distinct layers collaborate to effectively solve MMLU across all tested architectures. We now present evidence that these heads constitute part of a broader in-context retrieval mechanism. To establish this retrieval function, we analyze how removing one critical head impacts performance on a synthetic key-value (KV) retrieval task.
In the KV-Retrieval task, the model must recall a value associated with a given key from a predefined dictionary, testing its ability to store and access structured key-value pairs. For instance, given the prompt:
Memorize the following dictionary:
present:50
institute:0
scallops:84
neuropsychiatry:67
The value of the key ’scallops’ is
To succeed, the model must accurately store mappings (e.g., “scallops → 84”) and, when queried, retrieve the correct value (’84’) by identifying ’scallops’ in its stored representations without confusion.
We conducted experiments across 55 KV-Retrieval configurations with progressively increasing numbers of key-value pairs, each comprising 1,000 synthetic samples. For each configuration, we evaluated Llama-3.1-8B twice: once as-is and once with L17H24 (the head previously identified as critical in Section 3.4) removed. The results illustrated in Figure 4 reveal that removing L16H22 or L17H24 causes a decline in retrieval accuracy beyond just MMLU performance, confirming these heads’ essential role in the model’s retrieval mechanism. Notably, the impact of removing L17H24 grows with the number of key-value pairs, suggesting that its contribution to retrieval becomes more important as structural complexity increases. These findings extend previous research: the retrieval heads we identified align with and generalize the Content Gatherer head described by Lieberum et al. (2023) and the retrieval mechanisms documented by Wu et al. (2024). Our work demonstrates that similar retrieval mechanisms also emerge within SSM-based language models.
The pronounced role of retrieval in MMLU performance suggests that a broader Transformer-SSM performance gap may primarily stem from retrieval capabilities concentrated in just a few specific heads, rather than fundamental differences in language modeling capabilities. This insight fundamentally reframes our understanding of architectural tradeoffs, as the performance disparity appears to stem from a highly localized functional difference rather than a distributed architectural advantage.
Figure 4: Evaluation of Llama-3.1-8B on the KV-Retrieval task under three settings: unmodified, with only L16H22 removed, or with only L17H24 removed (the Gather and Aggregate Heads identified in Section 3.4, respectively). Performance is reported across 55 configurations with increasing numbers of key-value pairs, using the original (non-pruned) model. Each bar shows accuracy for a given dictionary size—for example, with 20 pairs, the original score is 85.6%, removing only the Gather Head (L16H22) drops accuracy to 76%, and removing only the Aggregate Head (L17H24) drops it to 62%. These results underscore two key points we build on later: (1) Retrieval is more sensitive to Aggregate Head removal than to Gather Head removal, underscoring its importance; (2) The Aggregate Head becomes increasingly important as task complexity grows.
In the next section, we will explore these two types of head and their distinct properties, revealing how they serve different roles.
3.6 Caveats
While the models we tested exhibit the same two-layer mechanism, the interaction does not always occur in consecutive layers. Instead, it can span non-adjacent layers, with intermediate steps in the residual stream. This separation reduces the effectiveness of pruning-based methods for identifying the interacting layers, even though the mechanism remains present. In such cases, alternative approaches, such as the activation patching used in Lieberum et al. (2023); Wiegreffe et al. (2025), may be more suitable to detect these interactions.
4 Gather-and-Aggregate Heads
In Section 3, we identified critical heads responsible for in-context retrieval across both Transformer- and SSM-based language models. We now examine a shared computational pattern underlying these heads: the Gather-and-Aggregate (G&A) mechanism.
The mechanism unfolds in two stages. First, a Gather Head isolates meaningful segments within the input—such as multiple-choice answers—and summarizes each into a representative token. Then, a later Aggregate Head uses these summaries to retrieve the information most relevant to the prompt.
While previous work has described retrieval-oriented heads in Transformers—such as the “Correct Letter” heads identified by Lieberum et al. (2023)—these analyses focused narrowly on multiple-choice tasks. Our findings show that such heads are part of a broader retrieval strategy that generalizes beyond this format. This perspective also helps refine the interpretation of results from Wu et al. (2024), who treated a group of retrieval heads as playing the same role. We find that these heads instead specialize in complementary functions within a coordinated process.
Finally, we show that this same mechanism emerges in SSM-based models, despite their lack of attention. These models exhibit smoother and less discrete patterns, but still learn to implement a G&A strategy using different architectural tools.
The remainder of this section unpacks the mechanism in detail. We begin with how it operates in Transformers (Sections 4.1 and 4.2), then present experimental evidence that it also underlies retrieval in SSMs (Section 4.3).
4.1 Gather Heads in Transformers
A Gather Head compresses semantically related tokens within a segment into a single representative token. In Transformer models, this typically occurs through attention patterns that cause the final token of each segment to “attend to” all prior tokens in the same segment, effectively summarizing them into a compact representation (see Figure 1(a)).
For example, consider the multiple-choice format introduced in Section 3:
[…]
While effective, this process is not always clean: attention sometimes leaks to unrelated tokens—particularly “attention sinks” Xiao et al. (2024) used to stabilize processing over long contexts. We return to these imperfections in Section 5.
4.2 Aggregate Heads in Transformers
Once the Gather Head condenses segments into representative tokens, an Aggregate Head—in a subsequent layer—retrieves and combines these summaries to inform the model’s prediction.
This head typically attends from a target token (e.g., the final “Answer:” position) over all gathered summaries, assigning greater weights to the most relevant one (see Figure 1(b)). This yields a sparse linear combination over segment summaries, often resembling an argmax-like operation (Tulchinskii et al., 2024):
[…]
In Transformer models, such heads were previously termed “Correct Letter” Heads (Lieberum et al., 2023), due to their consistent weighting of the correct answer’s summary token. However, the aggregation pattern generalizes beyond multiple-choice QA, enabling information retrieval in broader contexts.
As with Gather Heads, this mechanism is imperfect. Aggregate Heads sometimes distribute attention to irrelevant tokens, including attention sinks. These limitations become more pronounced in SSMs, as we discuss next.
4.3 G&A in Transformers and SSMs: Experimental Evidence
The G&A pattern is well-documented in Transformers. But do SSM-based models—like Mamba-1 and Mamba-2—implement the same structure? We answer this question through a two-part experimental procedure designed to isolate and test the role of specific heads:
  1. 1.
    Pruning: We prune each model to a minimal configuration (see Section 3.3), zeroing out all heads in the final two layers except those suspected to implement G&A (see Tables 6 and 1).
  2. 2.
    Masking: At runtime, we constrain the preserved heads’ attention patterns:
    • Gather Heads may only attend within each segment and to its final token.
    • Aggregate Heads may only attend to those final summary tokens.
    (See Figure 6 for an example.)
Across all models—Llama-3.1-8B (Transformer), Llamba-8B (Mamba-2), and Falcon-Mamba-7B (Mamba-1)—the masked and pruned models achieved identical MMLU accuracy as the unmodified versions.
This confirms that the selected heads are sufficient to drive retrieval and that the G&A mechanism is not unique to Transformers: despite differences in architecture, SSM-based models converge on the same two-stage strategy.
Nevertheless, the attention patterns differ in texture. As shown in Figure 5, attention maps in Mamba-based models are smoother and less sharply localized. This softness limits their ability to isolate segments cleanly:
  • Gather Heads in SSMs often leak attention outside segment boundaries.
  • Aggregate Heads struggle to sharply focus on the correct summary token.
These imperfections stem from the continuous nature of SSM hidden states, which blur boundary distinctions. We explore these architectural constraints further in Section 5.
Figure 5: Illustration of Gather Heads (top row) and Aggregate Heads (bottom row) across different model types. The left column shows attention patterns in Llama-3.1-8B (Transformer-based), while the center and right columns correspond to Llamba-8B (Mamba-2-based) and Falcon-Mamba-7B (Mamba-1-based), respectively—both state-space models (SSMs). In Transformer models, both Gather and Aggregate Heads produce sharp, localized patterns. In contrast, SSM-based models exhibit smoother behavior, making it harder to isolate specific segments or retrieve information with the same precision.
Figure 6: Visualization of the masking applied to the Gather Head (first row) and the Aggregate Head (second row) of Llamba-8B. The Gather Heads are restricted to interactions with the representative token and its associated answer, while the Aggregate Heads are limited to processing only the final token’s relevant portions. This setup ensures that the hypothesized Gather-and-Aggregate mechanism is exclusively responsible for the observed model behavior during MMLU evaluation.
5 The G&A Bottleneck in SSMs
While both empirical evidence and theoretical analysis have established that individual SSMs struggle with retrieval tasks (Jelassi et al., 2024; Wen et al., 2024), this limitation has largely been examined in full-scale language models as a black-box phenomenon (Jelassi et al., 2024; Arora et al., 2023; Waleffe et al., 2024). Here, we provide a mechanistic account of how SSMs’ retrieval difficulties manifest within language models, focusing on the Gather-and-Aggregate mechanism implemented by just a few critical heads. Since the Transformer–SSM gap is most evident in retrieval-heavy tasks, this analysis supports the view that the gap arises not from broad architectural deficiencies, but from a small number of heads that fail to implement key algorithmic operations.
This section presents a sequence of findings that build toward this conclusion:
  • Section 5.1 visually inspects Aggregate Heads in Transformer and SSM models, showing that SSMs struggle with sharp, selective aggregation—likely due to their hidden-state architecture.
  • Section 5.2 shows that SSMs require more Aggregate Heads to match Transformer performance, suggesting the network compensates for the limited expressivity of individual heads through duplication.
  • Section 5.3 demonstrates that hybrid models delegate aggregation to attention layers. Disabling just six attention heads causes a 4×4$\times$4 × drop in MMLU, indicating that SSMs alone cannot support strong retrieval.
  • Section 5.4 identifies the Aggregate Heads as the critical bottleneck: replacing a single layer that contains them boosts MMLU from 33% to 50%, confirming their central role in retrieval.
  • Section 5.5 outlines strategies to overcome this bottleneck in SSM-based models, suggesting that targeted attention—either added during training or preserved during distillation—can restore G&A functionality.
5.1 Visual Inspection of G&A Heads
As shown in Figures 5 and 6 and discussed in Section 4.2, the Aggregate Heads in Mamba-1 and Mamba-2 correctly attend to the final tokens of each segment. However, they also smoothly attend to neighboring tokens, even though Section 4.3 shows that they can function without them. This likely introduces noise into the linear combination of segment summaries, reducing aggregation effectiveness. This smooth behavior aligns with hidden-state architectures, where information is compressed into a single representation over time, making it harder to capture the global nature of Aggregate Heads.
5.2 Redundancy of Aggregate Heads in SSM Models
We find that SSM-based models require more Aggregate Heads than Transformer-based models to reach similar or even lower levels of performance. This suggests that SSMs compensate for limited per-head expressiveness by distributing the retrieval task across multiple weaker components. Three observations support this claim:
  1. 1.
    Gradual Degradation Reveals Redundant Retrieval Paths. As shown in Table 2 and discussed in Section 6, disabling heads in Llamba models leads to a slow, continuous drop in performance. This contrasts with the sharp collapses seen in Transformer-based models. The smoother degradation suggests that many SSM heads serve overlapping roles, indicating redundancy and less specialization.
  2. 2.
    Lack of Strong, Isolated Heads for Aggregation. Llamba-8B, distilled from Llama-3.1-8B using MOHAWK (Bick et al., 2024), mirrors the teacher’s structure by introducing G&A heads at layers 16 and 17. However, Table 7 shows a key difference: Llama-3.1-8B relies on two clearly effective heads (L17H24 and L17H27), while Llamba-8B depends on a single moderate head (L17H31), with others contributing minimally. Rather than isolating and concentrating aggregation in a few strong heads, Llamba spreads it across weaker alternatives.
  3. 3.
    More Heads Required to Match Performance. Llama-3.1-Minimal achieves 98% of its MMLU score with one Aggregate Head (Table 1). Llamba-8B-Minimal requires two to reach 95%, and Falcon-Mamba-7B-Minimal, built with lower-capacity Mamba-1 heads, needs four to reach 86% (Table 6). These differences indicate that SSM heads offer lower capacity individually and must be multiplied to maintain competitive performance.
5.3 Hybrid Models Delegate Aggregation to Attention
Attention layers in hybrid models are known to be critical for retrieval and copying tasks (Lieber et al., 2024; Glorioso et al., 2024; Waleffe et al., 2024). We hypothesize that during training, Aggregate Heads naturally concentrate in the attention layers, as attention’s greater expressivity makes it better suited for aggregation—underscoring a key limitation of SSMs.
To test this, we evaluate Zamba2-7B (Glorioso et al., 2024), the strongest 7B-scale hybrid model to date. Table 7 identifies attention heads critical for general retrieval; here, we further select nine that are especially important for MMLU: L47H{2, 4, 6, 8, 18, 21, 25} and L59H{17, 21}. Each head is disabled by zeroing out the final row of its attention matrix—a targeted intervention that impairs aggregation while leaving other capabilities, such as gathering and contextual mixing, largely intact. All SSM layers remain unchanged, preserving any aggregation they may contribute.
This intervention reduces MMLU accuracy from 64.3% to 34.9%, while knowledge task accuracy remains stable at 70%. Since MMLU’s random baseline is  25%, this 4x drop demonstrates that SSMs alone cannot sustain strong aggregation—hybrid models depend on attention layers to implement the G&A mechanism.
5.4 Hybrid Replacements
To further investigate whether SSMs struggle to implement the Aggregate Head, we examine how well an SSM mixer can replicate the functionality of an attention layer. We use the Llamba-8B (Bick et al., 2025) architecture, distilling each layer independently from Llama-3.1-8B (following Bick et al. (2024)) without end-to-end alignment. Specifically, the $i$-th Mamba-2 block in Llamba receives the output of Llama-3.1-8B’s $(i-1)$-th layer and is aligned independently using L2 distance against the output of Llama-3.1-8B’s $i$-th layer.
After assembling the full model from these individually distilled layers, we evaluate its performance on both knowledge tasks and MMLU. Interestingly, while the distilled model recovers 64% of Llama’s 69% on knowledge tasks, its MMLU performance remains 33%, compared to 67% for Llama. This suggests that some critical capability necessary for MMLU was not fully captured by the distilled layers.
To determine whether the observed performance gap is related to retrieval and specifically to the G&A mechanism, we conducted a hybrid replacement experiment. In this experiment, we systematically replaced each layer in the distilled Llamba-8B model with its corresponding layer from Llama-3.1-8B, evaluated its impact on MMLU (without fine-tuning), and then reverted the model before proceeding to the next layer.
The results, shown in Figure 7, reveal that most layer replacements had little or even a negative impact on MMLU. However, swapping L17, where we previously identified prominent Aggregate Heads at H24 and H27 (Table 1), significantly improved MMLU performance from 33% to 50%. To isolate their effect, we ran additional hybrid experiments: swapping only L17H24 while keeping the rest of the SSM heads intact increased performance to 44%, and including L17H27 raised it to 50%. This confirms that the contribution stems from the Aggregate Heads within L17, and further supports that individual SSM layers struggle to implement the Aggregate Head within G&A—reinforcing our claim that this limitation contributes to the performance gap between Transformers and SSMs.
5.5 Addressing G&A Limitations in SSM-Based Models
Efforts to improve the Gather-and-Aggregate mechanism within pure SSMs face a fundamental limitation: retrieval struggles stem from the fixed-size memory of RNN-style architectures, as shown in prior work (Jelassi et al., 2024; Wen et al., 2024). Rather than trying to overcome this constraint within the SSM design space, we explore hybrid models that incorporate attention layers to directly address these bottlenecks.
Our analysis reveals two empirical patterns that guide hybrid model design—one for training from scratch, and one for distillation:
  1. 1.
    Training hybrids from scratch. G&A heads consistently emerge in the middle of the network—for example, at layers L16 and L17 out of 32 in Llama-3.1-8B and Llamba-8B, and L35 and L36 out of 64 in Falcon-Mamba-8B. This suggests that attention is most valuable in mid layers, where retrieval operations naturally occur. As a result, when designing hybrid models from scratch, placing attention layers near the middle offers a targeted way to support G&A without relying on full attention throughout the network.
  2. 2.
    Distilling hybrids from a Transformer. As shown in Section 5.4, replacing individual layers in a distilled SSM with their attention-based counterparts from the teacher yields little benefit—except at L17, where MMLU improves from 33% to 50%. This indicates that the performance gap is concentrated in the Aggregate Head. Based on this, a practical strategy is to retain attention in the layers where strong G&A heads appear and distill the rest into SSMs. This allows the hybrid to recover retrieval quality where it matters most, while still benefiting from the efficiency of SSMs elsewhere.
These observations provide a practical recipe for hybrid design: introduce attention selectively—either guided by where G&A heads naturally emerge when training from scratch, or retained only in key locations when distilling from a Transformer. We leave further investigation into optimal placement and architectural variants of hybrid models to future work.
Figure 7: Results of the hybrid replacement experiment. Each Llamba-8B layer was replaced with its Llama-3.1-8B counterpart and tested on MMLU. Most swaps had little or negative impact on the 33% baseline, but replacing layer 17—where a key Aggregate Head is (Table 1)—significantly improved MMLU. This underscores both the difficulties SSMs face in replicating the Aggregate mechanism and G&A critical role in MMLU performance.
6 Transformer-SSM Gap Beyond MMLU
We previously showed that both Transformer- and SSM-based language models rely on a small number of heads to implement a two-stage Gather-and-Aggregate mechanism that supports in-context retrieval. This section extends that analysis to a broader range of benchmarks and model families, showing that differences in retrieval performance—and the gap between SSMs and Transformers in particular—often stem from specific algorithmic limitations, rather than from general deficiencies in language modeling ability.
We approach this in three stages: first, we isolate and characterize G&A heads across architectures; second, we quantify how architectural differences in G&A implementation affect task performance; and finally, we show how these differences are amplified or diminished by variations in task format.
6.1 Isolating Retrieval Differences Across Architectures
In Section 3, we showed that MMLU performance can be severely degraded by removing a single critical head—even when the model performs poorly on knowledge-focused tasks. This scenario is specific to MMLU: a task where strong retrieval alone can yield high accuracy, even when the model’s knowledge is degraded. In contrast, most benchmarks require both retrieval and factual knowledge, making it harder to isolate the contribution of retrieval alone.
As a result, the minimal-model strategy used in MMLU fails on other tasks: pruning layers diminishes knowledge capacity and lowers scores across the board, masking the effect of retrieval-specific mechanisms. To avoid this confound, we instead take a different approach: rather than pruning the model, we preserve its knowledge and selectively disable only the G&A heads.
As described in Section 3.5, the synthetic KV-Retrieval task allows us to identify heads involved in G&A by measuring the impact of ablating each head on retrieval accuracy (Table 7). A significant drop in performance indicates that the head plays a role in the G&A mechanism. A similar approach was also used by Wu et al. (2024) to identify retrieval heads in Transformers. Crucially, because KV-Retrieval is synthetic and does not require external knowledge, it isolates retrieval ability without interference from knowledge degradation. This lets us trace G&A heads across layers and model families in a controlled and generalizable way.
6.2 Quantifying the Retrieval Gap
Having identified the retrieval-critical heads, we now measure their practical impact across benchmarks. Specifically, we incrementally disable the identified G&A heads in three model families:
  • Transformers: Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct (Touvron et al., 2023), which serve as baselines for effective G&A implementation.
  • Pure SSMs: Llamba-3B and Llamba-8B (Bick et al., 2025), built exclusively with state-space layers based on Mamba-2.
  • Hybrids: Zamba2-2.7B and Zamba2-7B (Glorioso et al., 2024), architectures primarily based on SSM layers supplemented by shared attention modules.
Table 2 summarizes the results.
In Zamba2 hybrid models, we ablate only attention heads while leaving SSM layers intact. As shown in Section 5.3, these models primarily perform retrieval via the attention layers, making them the relevant target for G&A ablation. Additionally, because attention is shared across layers in Zamba2, we disable just one instance of each shared head at a time, treating it as non-shared to isolate its effect.
In Transformer models, we observe a steep decline in performance on retrieval-heavy tasks after disabling only 10–20 heads, while accuracy on knowledge-focused tasks remains stable. This confirms that retrieval depends on a small, specialized set of heads. Zamba2-7B exhibits a similar pattern, albeit with nuances due to its shared-attention structure. Notably, different occurrences of the same shared head behave similarly when ablated, consistent with shared G&A roles, as seen in Table 7.
In contrast, Llamba-8B degrades gradually. Even after removing 30 G&A heads, it retains over 95% of its original performance across both retrieval and knowledge tasks. This slow decline reflects redundancy: SSM-based models rely on many weaker heads to approximate the function of fewer, sharper Transformer heads—aligning with our observations from Section 5.
Model #Heads MMLU LAMB. GSM8K SWDE BBH Knowledge
acc $\uparrow$ ppl $\downarrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$
Llama-3B 0 60.3 (+0.0%) 4.8 (+0.0%) 28.7 (+0.0%) 85.8 (+0.0%) 38.2 (+0.0%) 60.5 (+0.0%)
10 53.1 (-12.0%) 6.5 (+35.7%) 17.4 (-39.4%) 81.9 (-4.5%) 33.4 (-12.6%) 59.4 (-1.8%)
20 32.2 (-46.6%) 8.8 (+82.8%) 9.1 (-68.2%) 57.5 (-33.0%) 27.7 (-27.5%) 58.7 (-3.0%)
30 29.9 (-50.4%) 10.1 (+109%) 5.6 (-80.5%) 47.5 (-44.6%) 25.4 (-33.5%) 58.0 (-4.1%)
Llama-8B 0 68.1 (+0.0%) 3.4 (+0.0%) 27.3 (+0.0%) 90.8 (+0.0%) 45.1 (+0.0%) 68.5 (+0.0%)
10 61.9 (-9.1%) 4.2 (+22.0%) 21.7 (-20.5%) 87.3 (-3.9%) 37.7 (-16.5%) 67.1 (-2.0%)
20 38.1 (-44.0%) 6.8 (+98.6%) 9.4 (-65.6%) 79.5 (-12.4%) 29.2 (-35.2%) 64.8 (-5.4%)
30 38.7 (-43.2%) 7.3 (+115%) 7.8 (-71.4%) 74.0 (-18.5%) 29.0 (-35.7%) 64.4 (-6.0%)
Llamba-3B 0 52.5 (+0.0%) 3.6 (+0.0%) - 21.3 (+0.0%) 9.2 (+0.0%) 63.8 (+0.0%)
10 42.6 (-18.9%) 5.2 (+44.4%) - 18.6 (-12.7%) 9.0 (-2.2%) 63.7 (-0.2%)
20 41.3 (-21.3%) 8.2 (+128%) - 18.1 (-15.0%) 9.0 (-2.2%) 63.1 (-1.1%)
30 41.2 (-21.5%) 9.1 (+153%) - 18.1 (-15.0%) 9.0 (-2.2%) 62.6 (-1.9%)
Llamba-8B 0 60.7 (+0.0%) 4.09 (+0.0%) - 20.0 (+0.0%) 11.0 (+0.0%) 69.1 (+0.0%)
10 59.6 (-1.8%) 6.26 (+53.2%) - 18.3 (-8.5%) 11.6 (+5.5%) 68.8 (-0.4%)
20 59.0 (-2.8%) 6.99 (+70.8%) - 18.6 (-7.0%) 11.4 (+3.6%) 68.6 (-0.7%)
30 58.1 (-4.3%) 7.44 (+81.9%) - 18.1 (-9.5%) 11.4 (+3.6%) 68.2 (-1.3%)
Zamba2-2.7B 0 55.7 (+0.0%) 4.2 (+0.0%) 57.4 (+0.0%) 89.5 (+0.0%) 30.6 (+0.0%) 66.8 (+0.0%)
10 42.4 (-23.9%) 12.8 (+204%) 24.7 (-57.0%) 84.3 (-5.8%) 25.5 (-16.7%) 64.8 (-3.0%)
20 37.2 (-33.2%) 22.2 (+428%) 6.5 (-88.7%) 74.4 (-16.9%) 17.4 (-43.1%) 62.6 (-6.3%)
Zamba2-7B 0 65.1 (+0.0%) 3.1 (+0.0%) 60.5 (+0.0%) 91.7 (+0.0%) 33.0 (+0.0%) 70.6 (+0.0%)
10 63.0 (-3.2%) 3.6 (+16.1%) 40.1 (-33.7%) 80.0 (-12.8%) 24.6 (-25.5%) 68.7 (-2.7%)
20 57.3 (-12.0%) 5.2 (+67.7%) 27.6 (-54.4%) 75.1 (-18.1%) 28.9 (-12.4%) 67.5 (-4.4%)
30 54.0 (-17.1%) 7.0 (+126%) 18.4 (-69.6%) 43.7 (-52.3%) 24.0 (-27.3%) 67.3 (-4.7%)
40 50.6 (-22.3%) 9.5 (+206%) 14.9 (-75.4%) 41.2 (-55.1%) 21.7 (-34.2%) 67.0 (-5.1%)
50 37.7 (-42.1%) 13.9 (+348%) 7.5 (-87.6%) 39.6 (-56.8%) 18.5 (-43.9%) 67.0 (-5.1%)
60 36.2 (-44.4%) 19.8 (+538%) 7.2 (-88.1%) 39.6 (-56.8%) 15.9 (-51.8%) 66.5 (-5.8%)
Table 2: Impact of Disabling G&A Heads Across Models. Each value reports the change relative to the 0-head configuration. Evaluations are conducted on MMLU (Hendrycks et al., 2021), LAMBADA (Paperno et al., 2016), Grade School Math (GSM8K) (Cobbe et al., 2021), SWDE (Arora et al., 2024), Big-Bench Hard (BBH) (Suzgun et al., 2022), and knowledge-focused tasks defined in Section 3.1. Note that (1) GSM8K and BBH are evaluated using the “exact-match, flexible-extract” metric; (2) GSM8K is omitted for Llamba models, which were not trained on mathematical datasets and are not intended for arithmetic reasoning.
6.3 Amplifying the Retrieval Gap through Task Format
Beyond architectural differences, the retrieval gap between Transformers and SSMs can be further amplified or mitigated by the task format itself. We previously saw hints of this in MMLU’s evaluation format, where rather than scoring the likelihood of full-text answer candidates, models must select a letter label (A, B, C, or D) corresponding to the correct option. To illustrate this effect, we compare two task variants with identical underlying content but different formats: ARC-Challenge in its original multiple-choice form versus ARC-Challenge Chat, a conversational variant.
As Table 3 shows, disabling G&A heads has minimal effect on the multiple-choice ARC-Challenge but drastically reduces accuracy in ARC-Challenge Chat. The conversational format thus increases retrieval demands by distributing relevant context throughout the dialogue, directly benefiting architectures with sharp retrieval mechanisms (Transformers and hybrids) over architectures with diffuse retrieval implementations (pure SSMs).
In this subsection, we examine a more subtle source of variation: the format of the task. Specifically, we compare two versions of ARC-Challenge that contain the same underlying questions but differ in presentation: the original multiple-choice format (used in our knowledge benchmarks), and a chat-style variant (ARC-Challenge Chat) shown in Table 3. While both versions rely on identical factual information, they respond very differently to G&A ablation.
ARC-Challenge Chat exhibits sharp performance drops—similar to MMLU—suggesting that its conversational structure introduces retrieval demands not present in the standard format. This highlights a key insight: retrieval load is shaped not only by what a task tests, but also by how it presents the information. In-context retrieval becomes more essential when relevant content is distributed across a dialogue, even if the underlying knowledge remains unchanged.
Model #Removed Heads ARC-Challenge (Chat) ARC-Challenge (Regular)
acc $\uparrow$ acc $\uparrow$
Llama-3B 0 76.8 (+0.0%) 45.5 (+0.0%)
10 72.2 (-6.0%) 43.6 (-4.2%)
20 50.0 (-34.9%) 42.0 (-7.7%)
30 43.2 (-43.8%) 41.9 (-7.9%)
Llama-8B 0 84.3 (+0.0%) 54.9 (+0.0%)
10 77.1 (-8.5%) 51.6 (-6.0%)
20 49.3 (-41.5%) 47.3 (-13.8%)
30 53.6 (-36.4%) 47.9 (-12.8%)
Table 3: Comparison of ARC Challenge and ARC-Challenge Chat under Attention Head Removal. ARC-Challenge Chat repackages the same questions as ARC Challenge into a conversational, generative format. While ARC Challenge uses normalized accuracy, the chat version relies on exact-match due to its open-ended outputs. Models initially perform better on the chat variant, leveraging its generative flexibility to reason through questions. But when G&A heads are ablated, performance drops sharply—unlike the minor declines seen in ARC Challenge. This contrast suggests the degradation reflects impaired retrieval rather than loss of stored knowledge.
7 Conclusions
In this work, we show how both Transformer and SSM models develop Gather-and-Aggregate (G&A) mechanisms to support in-context retrieval, which in turn shapes performance on many popular benchmarks. While SSMs struggle to implement effective Aggregate Heads due to their fixed-size hidden state, this does not render them ineffective for language modeling.
We find that end-to-end models rely on only a small number of G&A heads for retrieval, while most other heads—supporting abilities such as parametric knowledge extraction and language understanding—perform comparably across architectures. This suggests that the overall Transformer-SSM performance gap may stem from specific algorithmic limitations confined to a few heads.
Hybrid models provide further insight: they naturally delegate the G&A mechanism to attention layers, leveraging their stronger algorithmic capacity. This explains why even a small number of attention layers can significantly enhance SSM-based models, and offers practical guidance for designing future hybrid architectures.
Beyond improving our understanding of sequence model mechanisms, our findings also offer insights into popular benchmarks and the distinction between knowledge and algorithmic skills. While MMLU is often regarded as a gold-standard benchmark that requires broad knowledge, our results indicate that MMLU performance is primarily governed by retrieval ability. This is evident from our finding that a 4B model, despite underperforming on knowledge-focused benchmarks, can still achieve strong MMLU performance (66%). This suggests that algorithmic skills—particularly retrieval—rather than pure knowledge retention, play a dominant role in MMLU success.
While this work highlights the role of G&A in both attention and SSM models, important open questions remain. Narrowing the retrieval gap is a step forward, but in-context learning and copying mechanisms warrant further investigation, as they may also relate to G&A. Furthermore, our hybrid model analysis focused on a limited set of architectures; a broader study could refine the optimal balance of attention and SSM layers.
Appendix A Complementary Material for Section 3
In this section, we present additional results from the experiments conducted on the models explored in Section 3. These results provide further insights into the behavior of knowledge and retrieval capabilities across layers and heads in different models.
A.1 Detailed Performance Across Layers
We extend the analysis presented in Section 3.2 by providing performance data on other models, which show similar trends. As demonstrated in Figure 8, the pattern of stable MMLU scores despite a gradual decline in knowledge task performance is consistent across all models tested. Notably, the sharp drop in MMLU performance upon removal of specific layers reaffirms the criticality of certain layers in encoding retrieval skills.
Figure 8: Residual performance trends for knowledge tasks and the MMLU benchmark as layers are progressively pruned layers of Llama-3.1-8B, Llamba-8B, and Falcon-Mamba-7B. The leftmost bar represents the full model with all layers intact, while subsequent bars correspond to models with layers pruned incrementally from the end of the network. The plot highlights three key observations: (1) MMLU scores remain relatively stable despite a decline in general knowledge tasks performance, indicating reliance on task-specific skills; (2) a sharp drop in MMLU performance identifies a critical layer encoding essential skills; and (3) the gradual decline in knowledge tasks’ scores reflects the distributed nature of knowledge representation across layers.
A.2 Layer Removal: Mixer and MLP Components
For both Llama-3.1-8B and Llamba-8B, we performed experiments where entire layers, including both the mixer and MLP components, were removed. This approach was taken under the assumption that MLP components likely depend on outputs from the mixer components, which are crucial for information flow across layers. The results from these tests, presented in Figure 9, show that similar performance trends are observed whether the MLP components are retained or not. This suggests that the mixer components play a more prominent role in encoding retrieval capabilities, while the MLP components do not significantly alter the overall performance on knowledge tasks and MMLU benchmarks. This finding reinforces the idea that retrieval skills are primarily concentrated in specific layers and do not rely heavily on the MLP components within those layers.
Figure 9: Residual performance trends for knowledge tasks and the MMLU benchmark as layers are progressively pruned from Llama-3.1-8B, Llamba-8B but keeping the MLPs intact. This plot yields similar results to Figure 8
A.3 Performance in Minimal Models
The results from Section 3.3 are further elaborated in Figure 10. The performance of Llama-3.1-8B and other models when subjected to layer removal shows that, for all models, not one, but two layers are crucial for MMLU performance. The surprising robustness of MMLU performance in the minimal models highlights the importance of fine-tuning layer-wise architecture to achieve strong retrieval performance while retaining knowledge capabilities.
Figure 10: Residual performance trends for the MMLU benchmark and knowledge tasks metrics in the minimal model (defined with respect to MMLU) for Llama-3.1-8B, Llamba-8B, and Falcon-Mamba-7B. In this experiment, each layer is independently pruned, and the resulting performance is assessed. The results reveal that two layers are critical for MMLU performance, as pruning either of the last two layers causes a significant drop in MMLU scores while leaving knowledge metrics largely unaffected. This finding highlights the unique role of these layers in task-specific skills rather than general knowledge representation.
A.4 Critical Heads in MMLU Performance
As highlighted in Section 3.4, we identify specific heads within the critical layers that are responsible for maintaining high MMLU scores. Table 1 presents data on the performance of Llama-3.1-8B-Minimal when different configurations of heads are retained. These results underscore the significance of certain heads, where even a single mixing head can carry enough information to achieve a near state-of-the-art MMLU score. The removal of these heads leads to a drastic drop in MMLU performance, confirming their essential role in retrieval tasks.
These results are complemented by the data for Falcon-Mamba-8B and Llamba-8B in Table 6 and Table 5, which reveal similar trends of critical head importance across different model architectures.
Heads Kept in a Layer Metrics (%)
0-15 16 17 MMLU Knowledge Tasks
0,1,…,31 22 24 66.32 39.09
0,1,…,31 $\emptyset$ 24 24.36 39.18
0,1,…,31 22 $\emptyset$ 25.59 39.21
0,1,…,31 $\emptyset$ $\emptyset$ 25.56 39.21
Table 4: Evaluating the Role of Critical Heads in Llama-3.1-8B-Minimal’s Performance. This table presents the impact of two critical attention heads (L16H22 and L17H25) on MMLU performance while keeping all preceding layers intact. Four configurations were tested: (1) both heads retained, (2) only the last-layer head retained, (3) only the second-to-last-layer head retained, and (4) neither head retained. Results show that these heads must work together to sustain high MMLU performance ($>$66%), while removing either or both drops accuracy to random guessing (25%). Note: This is a duplicate of Table 1, reprinted here for convenience in comparison with Tables 6 and 5.
Heads Kept in a Layer Metrics (%)
0-15 16 17 MMLU Knowledge Tasks
0,1,…,31 10 3,9 58.92 39.08
0,1,…,31 $\emptyset$ 3,9 42.17 39.16
0,1,…,31 10 $\emptyset$ 38.02 39.0
0,1,…,31 no L17MLP $\emptyset$ 29.82 36.76
Table 5: Performance evaluation of Llamba-8B-Minimal, focusing on the impact of three critical heads identified in L16H10, L17H3, and L17H9. The experiment examines MMLU performance under four configurations: (1) no critical heads retained, (2) only the 22nd head retained, (3) only the 24th head retained, and (4) both heads retained. The findings demonstrate that the coordination of these two heads is crucial for maintaining high MMLU performance, achieving over 58%. Removing either head results in a significant performance decline of nearly 20%. Additionally, eliminating the MLP learned through these heads causes a further reduction of approximately 10%. This additional drop is likely attributable to the presence of another G&A mechanism in the shallower layers.
Heads Kept in a Layer Metrics (%)
0-34 35 36 MMLU Knowledge Tasks
0,…,8191 3186, 5143 6607, 7305 1565, 1906 3873, 4925 52.26 37.23
0,…,8191 $\emptyset$ 1565, 1906 3873, 4925 35.91 37.25
0,…,8191 3186, 5143 6607, 7305 $\emptyset$ 39.39 37.26
0,…,8191 $\emptyset$ $\emptyset$ 24.98 37.28
Table 6: Performance evaluation of Falcon-Mamba-8B-Minimal under various configurations involving critical mixing heads. The experiment isolates four specific heads identified in the last two layers and evaluates MMLU scores in the following configurations: (1) all critical heads removed, (2) critical heads from the first layer retained, (3) critical heads from the second layer retained, and (4) all critical heads retained. The results reveal that only four heads, out of thousands at every layer are essential for leveraging the network’s knowledge to achieve an MMLU score of 52.26%. Excluding these heads results in a sharp performance drop to 25%, illustrating their critical role in transforming distributed knowledge into task-specific capabilities.
A.5 Retrieval Heads Across Model Families
Table 7 presents the 50 most retrieval-relevant heads for each model, ranked by their individual contribution to KV-Retrieval accuracy. Several trends emerge. In Transformers, top heads cluster in upper layers, often overlapping with those critical for MMLU. In SSM-based models, retrieval heads are more diffuse and individually weaker. Zamba2 hybrids concentrate high-performing heads around shared-attention layers, suggesting these layers anchor retrieval in otherwise SSM-heavy architectures.
Table 7: Attention Head Ablation Study: Performance impact on KV-Retrieval task when individual attention heads are removed from various language models. Each head was systematically knocked out and the model’s accuracy was evaluated on a KV-Retrieval benchmark with dictionary size of 20 across 1000 samples. After measurement, each head was restored before testing the next one. Results show the 50 most significant attention heads from each model, sorted by accuracy, revealing which heads contribute most critically to the KV-Retrieval capability across Zamba2, Llama, and Llamba model families. For Zamba2 models, marks layers with the first shared attention, and marks those with the second.
Zamba2-2.7B Zamba2-7B Llama-3.2-3B Llama-3.1-8B Llamba-3B Llamba-8B
Layer Head Acc Layer Head Acc Layer Head Acc Layer Head Acc Layer Head Acc Layer Head Acc
6 0 0.088 11 28 0.139 14 7 0.381 12 1 0.421 20 30 0.214 17 31 0.382
18 12 0.199 23 28 0.139 12 1 0.405 8 8 0.494 1 0 0.220 21 2 0.446
6 12 0.303 35 28 0.139 14 6 0.524 13 5 0.514 15 30 0.296 16 29 0.460
30 0 0.328 47 28 0.139 15 18 0.545 14 7 0.519 14 17 0.301 22 26 0.465
6 26 0.380 59 28 0.176 13 5 0.589 13 17 0.596 11 23 0.317 27 15 0.471
30 4 0.421 71 28 0.176 10 17 0.597 15 18 0.635 12 24 0.320 14 9 0.471
18 25 0.426 11 21 0.266 9 1 0.625 14 6 0.680 19 13 0.324 11 25 0.476
30 20 0.430 23 21 0.266 0 4 0.639 13 23 0.685 19 7 0.327 21 27 0.482
18 26 0.433 35 21 0.266 12 13 0.648 12 2 0.730 10 20 0.329 16 18 0.487
30 31 0.472 47 21 0.266 9 18 0.655 5 14 0.732 15 6 0.333 11 31 0.487
24 23 0.479 59 21 0.281 12 8 0.657 10 17 0.761 24 17 0.334 10 15 0.490
30 12 0.482 71 21 0.281 14 0 0.667 11 9 0.784 13 29 0.335 14 5 0.492
30 24 0.483 11 3 0.424 15 20 0.670 5 22 0.816 7 28 0.337 11 29 0.492
30 26 0.488 23 3 0.424 10 3 0.671 14 0 0.828 10 16 0.338 15 16 0.495
36 3 0.496 35 3 0.424 3 18 0.677 19 7 0.830 14 6 0.340 12 9 0.495
36 30 0.496 47 3 0.424 19 7 0.677 9 23 0.839 12 25 0.344 18 30 0.497
30 16 0.499 6 2 0.430 8 8 0.685 7 16 0.844 11 16 0.348 14 7 0.497
18 16 0.500 17 2 0.430 11 10 0.689 12 23 0.846 14 9 0.350 28 28 0.498
6 16 0.508 29 2 0.430 12 0 0.689 10 11 0.855 19 0 0.352 22 28 0.498
24 3 0.514 41 2 0.430 13 16 0.693 16 15 0.855 13 19 0.352 9 28 0.501
24 6 0.521 53 2 0.430 9 21 0.696 11 11 0.859 10 17 0.352 22 9 0.502
24 7 0.521 65 2 0.434 11 1 0.698 7 4 0.860 16 22 0.353 9 20 0.502
30 25 0.524 77 2 0.434 12 12 0.699 12 7 0.868 12 10 0.353 11 0 0.504
24 21 0.526 11 15 0.473 3 2 0.703 10 20 0.870 2 9 0.355 8 17 0.504
18 20 0.532 11 18 0.473 1 17 0.705 3 4 0.878 0 7 0.355 15 9 0.505
6 1 0.532 23 15 0.473 5 13 0.705 0 7 0.879 1 23 0.356 6 10 0.505
24 25 0.533 23 18 0.473 13 22 0.705 4 5 0.883 8 26 0.357 5 24 0.505
12 12 0.534 35 15 0.473 9 3 0.710 15 20 0.883 0 18 0.357 19 24 0.506
24 22 0.538 35 18 0.473 6 4 0.713 1 7 0.884 11 24 0.359 15 17 0.506
36 16 0.539 47 15 0.473 10 18 0.713 3 23 0.885 16 9 0.360 12 18 0.506
24 13 0.539 47 18 0.473 5 19 0.714 10 10 0.886 13 30 0.360 7 11 0.507
18 10 0.541 59 3 0.480 12 10 0.717 12 3 0.886 11 0 0.352 24 5 0.508
30 6 0.544 71 3 0.480 11 15 0.721 12 6 0.886 13 19 0.352 13 29 0.508
18 21 0.545 11 17 0.488 9 17 0.722 9 11 0.888 7 7 0.360 7 20 0.508
12 19 0.545 23 17 0.488 4 4 0.726 11 1 0.890 6 1 0.360 7 30 0.508
18 19 0.546 35 17 0.488 11 2 0.729 8 17 0.892 26 6 0.361 6 18 0.508
24 15 0.547 47 17 0.488 12 4 0.730 1 15 0.895 23 16 0.361 5 8 0.508
18 24 0.548 59 15 0.488 10 8 0.732 4 4 0.896 19 9 0.361 13 31 0.509
6 25 0.548 71 15 0.488 13 21 0.733 6 15 0.896 7 9 0.361 12 15 0.509
18 0 0.549 11 10 0.496 2 16 0.735 9 18 0.896 22 27 0.362 11 12 0.509
6 11 0.549 23 10 0.496 10 5 0.735 8 12 0.897 16 16 0.362 10 1 0.509
24 12 0.550 35 10 0.496 11 14 0.735 24 15 0.897 11 8 0.362 10 8 0.509
6 24 0.550 47 10 0.496 11 9 0.739 3 3 0.898 9 16 0.362 10 17 0.509
18 2 0.551 59 18 0.496 5 17 0.740 0 1 0.900 3 27 0.362 8 30 0.509
24 1 0.552 71 18 0.496 13 12 0.741 5 6 0.900 11 22 0.363 7 24 0.509
12 9 0.552 6 1 0.500 9 8 0.743 20 11 0.900 6 24 0.363 5 7 0.509
6 22 0.553 17 1 0.500 7 3 0.744 3 16 0.901 18 7 0.364 12 7 0.510
18 6 0.554 29 1 0.500 12 22 0.745 10 4 0.902 16 15 0.364 11 8 0.510
12 7 0.554 41 1 0.500 19 1 0.747 1 0 0.903 14 27 0.364 9 0 0.510
30 2 0.555 53 1 0.500 4 8 0.748 3 1 0.905 13 24 0.364 22 31 0.511